OpenStudy (anonymous):

The polar graph of the equation \[r = \frac{1}{2 - \cos\theta}\] has what shape? (Note: THis is supposedly Algebra II, so it's probably very basic polar. Therefore, I'm not very familiar with with polar coordinates and graphs.)

OpenStudy (jamesj):

Plot out what happens when theta = 0, pi/2, pi, 3pi/2. What does it look like?

OpenStudy (jamesj):

Once you have an intuitive feel for what this is, you can work out the algebra, by substituting r = sqrt(x^2+y^2), cos theta = x/r

OpenStudy (anonymous):

Well, isn't it a repeating parabola?

OpenStudy (jamesj):

No.

OpenStudy (jamesj):

for theta = 0, r = 1 theta = pi/2, r = 1/2 theta = pi, r = 1/3 theta = 3pi/2,r = 1/2 Sketch out those four points on the plane.

OpenStudy (anonymous):

I'm new to polar, so I had to get something to graph it for me. Isn't there a special graphing system for polar coordinates?

OpenStudy (anonymous):

Can you link me to something that's very basic (something that can teach me very basic polar) because I doubt I'll need to know much until Pre-Calculus

OpenStudy (jamesj):

Seriously begin with pen and paper with the four points above. What are the location of those four points in Cartesian coordinates? theta = 0, r = 1 Ths point lies along the direction of the positive x-axis, 1 from the origin. Hence its Cartesian coordinates are (1,0). Now you figure out the Cartesian coordinates for these three points: theta = pi/2, r = 1/2 theta = pi, r = 1/3 theta = 3pi/2,r = 1/2 ...and then plot all four points.

OpenStudy (anonymous):

THis is the sort of graph I keep on getting.

OpenStudy (jamesj):

No. What are the coordinates of theta = pi/2, r = 1/2? This is in the positive y direction. Hence another point is (0, 1/2) So two points are (1,0) and (0,1/2) What are the other two?

OpenStudy (anonymous):

Is another point (0, 1/3)? I'm seeing if I'm on the right track.

OpenStudy (jamesj):

theta = pi is in the *negative* y direction. So (0, 1/3) isn't quite right.

OpenStudy (anonymous):

So (0, -1/3)

OpenStudy (jamesj):

**correction, the negative x direction.

OpenStudy (jamesj):

theta = pi is the negative x direction, hence the third point is at (-1/3, 0) Now where's the fourth point?

OpenStudy (jamesj):

theta r (x,y) 0 1 (1, 0) pi/2 1/2 (0, 1/2) pi 1/3 (-1/3, 0) 3pi/2 1/2 ....

OpenStudy (anonymous):

(0, -1/2 ) Is this right?

OpenStudy (jamesj):

Yes. Now, plot those four points and tell me what kind of geometric figure would go through all four of them.

OpenStudy (anonymous):

I got an ellipse? These are the points I graphed.

OpenStudy (jamesj):

Yes, an ellipse.

OpenStudy (anonymous):

polar coordinates to rectangular coordinates \[x=r \cos \theta \\ r=\sqrt{x^2+y^2} \\ you \ \ have \ \ r=\frac{1}{2-\cos \theta} \\ 2r-r \cos \theta=1 \ \ then \ \ 2\sqrt{x^2+y^2}-x=1 \\ 2\sqrt{x^2+y^2}=x+1 \\4(x^2+y^2)=(x+1)^2\\ 4y^2=-3x^2+2x+1\\4y^2=-3(x^2-\frac{2}{3}x)+1\\4y^2=-3(x^2-\frac{2}{3}x+\frac{1}{9}-\frac{1}{9})+1\\4y^2=-3(x^2-\frac{2}{3}x+\frac{1}{9})+\frac{1}{3}+1\\4y^2+3(x-\frac{1}{3})^2=\frac{4}{3} \\ 3y^2+\frac{9}{4}(x-\frac{1}{3})^2=1\\ its \ \ an \ \ Ellipse \]

OpenStudy (anonymous):

Alright. Thanks! I didn't know how to convert polar coordinates to cartesian coordinates, so I just used the Unit Circle for this. Am I supposed to do that or is there a different process?

OpenStudy (jamesj):

What is the definition of polar coordinates?

OpenStudy (anonymous):

Isn't it a coordinate that represents the distance from a fixed point?

OpenStudy (jamesj):

Not quite. You should look this up in your text book and make sure you really understand it.

OpenStudy (anonymous):

Well, I don't have a text book because this is for an Algebra II math team test. I'm guessing to challenge people they added this question in because I don't think I'm supposed to know much about polar coordinates and are giving props to those that do for some reason. Anyway, thank you!

OpenStudy (jamesj):

Polar coordinates are defined as the distance from the origin, that is the variable r. And angle to the positive x axis in the anticlockwise direction, this is the variable theta.

OpenStudy (jamesj):

Draw yourself a diagram and convince yourself that x = r cos(theta) y = r sin(theta)

OpenStudy (anonymous):

I'm finding this very comparable to the unit circle except r isn't always 1.

OpenStudy (jamesj):

From this it follows that \[ x^2 + y^2 = r^2 (\cos^2 \theta + \sin^2 \theta) = r^2 \] i.e., \[ r =\sqrt{x^2 + y^2}\]

OpenStudy (anonymous):

Ok. Thank you!

OpenStudy (jamesj):

Also \[ \frac{y}{x} = \frac{r \sin \theta}{r \cos \theta} = \tan \theta \] Hence \[ \theta = \arctan(y/x) \]